OpenStudy (anonymous):

In the equation x/4 + x = 6, x = 1.2.

OpenStudy (anonymous):

true or false

OpenStudy (xapproachesinfinity):

adding up 1/4x+x=5/4x so 5/4x=6 multiply both side by 4/5 \[\frac{ 4 }{ 5 }\times \frac{ 5 }{ 4 }x=\frac{ 4 }{ 5}\times6\]

OpenStudy (xapproachesinfinity):

x=\fraction{24}{5}

OpenStudy (xapproachesinfinity):

\[x=\frac{ 24 }{ 5 }\]

OpenStudy (xapproachesinfinity):

So it 4.8 not 1.2 right?

OpenStudy (anonymous):

so it would be false

OpenStudy (xapproachesinfinity):

Correct!

OpenStudy (xapproachesinfinity):

Did you understand the steps?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

Good!